# Multidimensional empirical wavelet transform

### v1

@@ 3.3 Empirical wavelet transform
In this section, we introduce continuous and discrete transforms of a function $f\in{\mathrm{L}^{2}}({\mathbb{R}}^{N})$ based on either the empirical wavelet systems $\{\psi_{n}\}_{n\in\Upsilon}$ or the symmetric empirical wavelet systems $\{\phi_{n}\}_{n\in\Upsilon}$.
@@
${\mathcal{E}}_{\phi}^{f}(b,n)=\langle f,T_{b}\phi_{n}\rangle.$ (5)
In this definition, the $N$-dimensional continuous symmetric empirical wavelet transform ${\mathcal{E}}_{\phi}^{f}$ is symmetric for the frequency variable $n$ but not for the translation variable $b$.
@@
${\mathcal{E}}_{\phi}^{f}(b,n)=\left(f\star\overline{\phi^{*}_{n}}\right)(b)={\mathcal{F}}^{-1}\left(\widehat{f}\bullet\overline{\widehat{\phi}_{n}}\right)(b).$ (7)
@@
Given the relation between $\phi_{n}$ and $\psi_{n}$ in the spatial domain, which is given by Equation 3, Equation 7 directly stems from Equation 6 using the linearity of the convolution, the inner product and the inverse Fourier transform. ∎
@@
Let $\{b_{n}\}_{n\in\Upsilon}\in{\mathbb{R}}^{N}\setminus\{0\}$. The $N$-dimensional discrete empirical wavelet transform of a real or complex-valued function $f\in{\mathrm{L}^{2}}({\mathbb{R}}^{N})$ is defined by, for every $n\in\Upsilon$ and $k\in\mathbb{Z}^{N}$,
@@
${\mathcal{E}}_{\phi}^{f}(b_{n}k,n)=\langle f,T_{b_{n}k}\phi_{n}\rangle,$
@@ 4.1 Continuous frames
In this section, we build dual frames of the two proposed systems
$\{T_{b}\psi_{n}\}_{(n,k)\in\Upsilon\times{\mathbb{R}}^{N}}\quad\textrm{ and }\quad\{T_{b}\phi_{n}\}_{(n,k)\in\Upsilon^{+}\times{\mathbb{R}}^{N}},$
involved in the continuous wavelet transform of Definition 5. This allows to find sufficients conditions for these systems to be tight frames.
@@
$\sum_{n\in\Upsilon}\left({\mathcal{E}}_{\psi}^{f}(\cdot,n)\star\widetilde{\psi}_{n}\right)(x)$ $=\sum_{n\in\Upsilon}\int_{b\in{\mathbb{R}}^{N}}{\mathcal{E}}_{\psi}^{f}(b,n)\widehat{\psi}_{n}(x-b)\mathrm{d}b$
$=\sum_{n\in\Upsilon}\int_{b\in{\mathbb{R}}^{N}}\langle f,T_{b}\psi_{n}\rangle T_{b}\widehat{\psi}_{n}(x)\mathrm{d}b.$
@@
$f(x)=\frac{1}{A}\sum_{n\in\Upsilon}\left({\mathcal{E}}_{\psi}^{f}(\cdot,n)\star\psi_{n}\right)(x).$
Then, we can rewrite
$f(x)$ $=\frac{1}{A}\sum_{n\in\Upsilon}\int_{b\in{\mathbb{R}}^{N}}{\mathcal{E}}_{\psi}^{f}(b,n)\psi_{n}(x-b)\mathrm{d}b$
$=\frac{1}{A}\sum_{n\in\Upsilon}\int_{b\in{\mathbb{R}}^{N}}\langle f,T_{b}\psi_{n}\rangle T_{b}\psi_{n}(x)\mathrm{d}b.$
Now, we consider the symmetric wavelet filterbank $\{\phi_{n}\}_{n\in\Upsilon^{+}}$. The following theorem guarantees the exact reconstruction of a real-valued function $f$ from the continuous symmetric empirical wavelet transform ${\mathcal{E}}_{\phi}^{f}$ given in Equation 5.
@@
$f(x)=\sum_{n\in\Upsilon^{+}}\left({\mathcal{E}}_{\phi}^{f}(\cdot,n)\star\tilde{\phi}_{n}\right)(x)=\sum_{n\in\Upsilon^{+}}\int_{b\in{\mathbb{R}}^{N}}\langle f,T_{b}\phi_{n}\rangle T_{b}\widetilde{\phi}_{n}(x)\mathrm{d}b,,$ (9)
where the set of dual symmetric empirical wavelets $\{\widetilde{\phi}_{n}\}_{n\in\Upsilon}$ is defined by, for every $n\in\Upsilon$ and $\xi\in{\mathbb{R}}^{N}$,
$\widehat{\widetilde{\phi}_{n}}(\xi)=\frac{\widehat{\phi}_{n}(\xi)}{\displaystyle\sum_{n\in\Upsilon^{+}}\left|\widehat{\phi}_{n}(\xi)\right|^{2}}.$
@@
$\sum_{n\in\Upsilon^{+}}\left|\widehat{\phi}_{n}(\xi)\right|^{2}$ $=\left|\widehat{\psi}_{0}(\xi)\right|^{2}+\frac{1}{2}\sum_{n\in\Upsilon^{+}}\left|\widehat{\psi}_{n}(\xi)+\widehat{\psi}_{-n}(\xi)\right|^{2}$
@@
which ensures that the symmetric empirical wavelet filters $\widetilde{\phi}_{n}$ are well defined. Then, Equation 9 stems from the same computation as in the proof of Theorem 1. ∎
@@
then $\{T_{b}\phi_{n}\}_{(n,b)\in\Upsilon^{+}\times{\mathbb{R}}^{N}}$ is a continuous tight frame. Specifically, for every $x\in{\mathbb{R}}^{N}$,
$f(x)=\frac{1}{A+B}\sum_{n\in\Upsilon^{+}}\int_{b\in{\mathbb{R}}^{N}}\langle f,T_{b}\phi_{n}\rangle T_{b}\phi_{n}(x)\mathrm{d}b.$
@@
$\sum_{n\in\Upsilon^{+}}\left|\widehat{\phi}_{n}(\xi)\right|^{2}$ $=\frac{1}{2}\sum_{n\in\Upsilon^{+}}\left|\widehat{\psi}_{n}(\xi)\right|^{2}+\frac{1}{2}\sum_{n\in\Upsilon^{+}}\left|\widehat{\psi}_{-n}(\xi)\right|^{2}+\sum_{n\in\Upsilon^{+}\setminus\{0\}}\left|\widehat{\psi}_{n}(\xi)\right|\left|\widehat{\psi}_{-n}(\xi)\right|=A+B,$
and the result follows from a computation similar to the one of the proof of Corollary 1. ∎
@@ 4.2 Discrete frames
In this section, we exhibit conditions to build discrete wavelet frames involved in the wavelet transforms of Definition 6. These conditions depend on the compactness of the support $\overline{\Lambda}$ of the Fourier transform of the mother wavelet $\widehat{\psi}$. The two underlying cases are examined separately.
@@ 4.2.1 Compactly supported $\boldsymbol{\widehat{\psi}}$
In this section, we consider that $\widehat{\psi}$ has a compact support $\overline{\Lambda}$. Excluding the supports of a partition $\{\Omega_{n}\}_{n\in\Upsilon}$ with non-compact closure, we can state sufficient and necessary conditions for which the systems
$\{T_{b_{n}k}\psi_{n}\}_{(n,k)\in\Upsilon_{comp}\times{\mathbb{Z}}^{N}}\quad\textrm{ and }\quad\{T_{b_{n}k}\phi_{n}\}_{(n,k)\in\Upsilon_{comp}^{+}\times{\mathbb{Z}}^{N}},$
@@
$\Upsilon_{comp}=\{n\in\Upsilon\mid\overline{\Omega_{n}}\textrm{ is compact}\}\;\textrm{ and }\;\Upsilon_{comp}^{+}=\{n\in\Upsilon^{+}\mid\overline{\Omega_{n}}\textrm{ is compact}\}.$
The following theorem first gives a sufficient and necessary condition to build a tight frame from $\{\psi_{n}\}_{n\in\Upsilon_{comp}}$. It is a straight generalization of Theorems 4-7 of [9] to the $N$-dimensional case.
@@
Let us denote $\mathrm{L}_{comp}^{2}({\mathbb{R}}^{N})=\{f\in\mathrm{L}^{2}({\mathbb{R}}^{N})\mid\mathrm{supp}\,\widehat{f}\subseteq\Gamma_{comp}\}$ and $\Gamma_{comp}=\underset{n\in\Upsilon_{comp}}{\bigcup}\overline{\Omega_{n}}$. The system $\{T_{b_{n}k}\psi_{n}\}_{(n,k)\in\Upsilon\times{\mathbb{Z}}^{N}}$ is a Paraseval frame for $\mathrm{L}_{comp}^{2}({\mathbb{R}}^{N})$ if and only if, for $a.e.\,\xi\in{\mathbb{R}}^{N}$,
@@
$\mathcal{K}=\underset{n\in\Upsilon_{comp}}{\bigcup}b_{n}^{-1}\mathbb{Z}^{N},\qquad\Upsilon_{\alpha}=\{n\in\Upsilon_{comp}\mid b_{n}\alpha\in\mathbb{Z}^{N}\},$
@@
Let us denote $\mathcal{D}=\{f\in\mathrm{L}^{2}({\mathbb{R}}^{N})\mid\widehat{f}\in\mathrm{L}^{\infty}({\mathbb{R}}^{N})\textrm{ and }\mathrm{supp}\,\widehat{f}\subset\Gamma_{comp}\}.$ Theorem 2.1 in [12] states, with $g_{p}=\psi_{n}$, $C_{p}=b_{n}$ and $\mathcal{P}=\Upsilon_{comp}$, the desired equivalence under the condition that
$\forall f\in\mathcal{D},\;\sum_{n\in\Upsilon_{comp}}\sum_{k\in\mathbb{Z}^{N}}\int_{\mathrm{supp}\,\widehat{f}}\left|\widehat{f}(\xi+b_{n}^{-1}k)\right|^{2}\frac{1}{|b_{n}|}\left|\widehat{\psi}_{n}(\xi)\right|^{2}\mathrm{d}\xi<\infty.$
@@
Similarly, the following theorem gives a sufficient and necessary condition to build a tight frame from the symmetric filterbank $\{\phi_{n}\}_{n\in\Upsilon_{comp}^{+}}$.
Theorem 4 (Discrete continuous Parseval frame).
Let us denote $\mathrm{L}_{comp}^{2}({\mathbb{R}}^{N})=\{f\in\mathrm{L}^{2}({\mathbb{R}}^{N})\mid\mathrm{supp}\,\widehat{f}\subseteq\Gamma_{comp}\}$ and $\Gamma_{comp}=\underset{n\in\Upsilon_{comp}}{\bigcup}\overline{\Omega_{n}}$. Then, the system $\{T_{b_{n}k}\phi_{n}\}_{(n,b)\in\Upsilon^{+}\times{\mathbb{Z}}^{N}}$ is a Parseval frame for $\mathrm{L}_{comp}^{2}({\mathbb{R}}^{N})$ if and only if, for $a.e.\,\xi\in{\mathbb{R}}^{N}$,
$\sum_{n\in\Upsilon_{\alpha}^{+}}\frac{1}{|b_{n}|}\widehat{\phi}_{n}(\xi)\overline{\widehat{\phi}_{n}(\xi+\alpha)}=\delta_{\alpha,0},$ (10)
@@
$\mathcal{K}^{+}=\underset{n\in\Upsilon_{comp}^{+}}{\bigcup}b_{n}^{-1}\mathbb{Z}^{N},\qquad\Upsilon_{\alpha}^{+}=\{n\in\Upsilon_{comp}^{+}\mid b_{n}\alpha\in\mathbb{Z}^{N}\},$
@@
Let us denote $\mathcal{D}=\{f\in\mathrm{L}^{2}({\mathbb{R}}^{N})\mid\widehat{f}\in\mathrm{L}^{\infty}({\mathbb{R}}^{N})\textrm{ and }\mathrm{supp}\,\widehat{f}\subset\Gamma_{comp}\}.$ Noticing that
$\Gamma_{comp}=\underset{n\in\Upsilon_{comp}^{+}}{\bigcup}\left(\overline{\Omega_{n}}\cup\overline{\Omega_{-n}}\right).$
Theorem 2.1 in [12] states (by taking $g_{p}=\phi_{n}$, $C_{p}=b_{n}$ and $\mathcal{P}=\Upsilon_{comp}^{+}$) that the system $\{T_{b_{n}k}\phi_{n}\}_{(n,k)\in\Upsilon^{+}\times{\mathbb{Z}}^{N}}$ is a Paraseval frame for $\mathrm{L}_{comp}^{2}({\mathbb{R}}^{N})$ if and only if, for $a.e.\,\xi\in{\mathbb{R}}^{N}$,
$\sum_{n\in\Upsilon_{\alpha}^{+}}\frac{1}{|b_{n}|}\widehat{\phi}_{n}(\xi)\overline{\widehat{\phi}_{n}(\xi+\alpha)}=\delta_{\alpha,0},$
@@
$\forall f\in\mathcal{D},\;\sum_{n\in\Upsilon_{comp}^{+}}\sum_{k\in\mathbb{Z}^{N}}\int_{\mathrm{supp}\,\widehat{f}}\left|\widehat{f}(\xi+b_{n}^{-1}k)\right|^{2}\frac{1}{|b_{n}|}\left|\widehat{\phi}_{n}(\xi)\right|^{2}\mathrm{d}\xi<\infty.$
This condition is given by the proof of Theorem 4 in [9], replacing $\mathbb{Z}$ by $\mathbb{Z}^{N}$ and $\psi_{n}$ by $\phi_{n}$. It results from the fact that $\mathrm{supp}\,\widehat{f}$ and $\mathrm{supp}\,\widehat{\phi}_{n}$ are compact and that there are finitely many $\mathrm{supp}\,\widehat{\psi}_{n}$ that intersect $\mathrm{supp}\,\widehat{f}$. This proves the equivalence. ∎

### v2

@@ 3.3 Empirical wavelet transform
In this section, we introduce continuous and discrete transforms of a function $f\in{\mathrm{L}^{2}}({\mathbb{R}}^{N})$ based on either the empirical wavelet systems $\{\psi_{n}\}_{n\in\Upsilon}$ or the symmetric empirical wavelet systems $\{\chi_{n}\}_{n\in\Upsilon}$.
@@
${\mathcal{E}}_{\chi}^{f}(b,n)=\langle f,T_{b}\chi_{n}\rangle.$ (5)
In this definition, the $N$-dimensional continuous symmetric empirical wavelet transform ${\mathcal{E}}_{\chi}^{f}$ is symmetric for the frequency variable $n$ but not for the translation variable $b$.
@@
${\mathcal{E}}_{\chi}^{f}(b,n)=\left(f\star\overline{\chi^{*}_{n}}\right)(b)={\mathcal{F}}^{-1}\left(\widehat{f}\bullet\overline{\widehat{\chi}_{n}}\right)(b).$ (7)
@@
Given the relation between $\chi_{n}$ and $\psi_{n}$ in the spatial domain, which is given by Equation 3, Equation 7 directly stems from Equation 6 using the linearity of the convolution, the inner product and the inverse Fourier transform. ∎
@@
Let $\{b_{n}\}_{n\in\Upsilon}\in{\mathbb{R}}\setminus\{0\}$. The $N$-dimensional discrete empirical wavelet transform of a real or complex-valued function $f\in{\mathrm{L}^{2}}({\mathbb{R}}^{N})$ is defined by, for every $n\in\Upsilon$ and $k\in\mathbb{Z}^{N}$,
@@
${\mathcal{E}}_{\chi}^{f}(b_{n}k,n)=\langle f,T_{b_{n}k}\chi_{n}\rangle,$
@@ 4.1 Continuous frames
In this section, we build dual frames for the two proposed systems
$\{T_{b}\psi_{n}\}_{(n,b)\in\Upsilon\times{\mathbb{R}}^{N}}\quad\textrm{ and }\quad\{T_{b}\chi_{n}\}_{(n,b)\in\Upsilon^{+}\times{\mathbb{R}}^{N}},$
with $\Upsilon^{+}=\{n\in\Upsilon\mid n\geq 0\}$, involved in the continuous wavelet transform of Definition 5. This allows to find sufficients conditions for these systems to be tight frames.
@@
$\sum_{n\in\Upsilon}\left({\mathcal{E}}_{\psi}^{f}(\cdot,n)\star\widetilde{\psi}_{n}\right)(x)$ $=\sum_{n\in\Upsilon}\int_{b\in{\mathbb{R}}^{N}}{\mathcal{E}}_{\psi}^{f}(b,n)\widetilde{\psi}_{n}(x-b)\mathrm{d}b$
$=\sum_{n\in\Upsilon}\int_{b\in{\mathbb{R}}^{N}}\langle f,T_{b}\psi_{n}\rangle T_{b}\widetilde{\psi}_{n}(x)\mathrm{d}b.$
@@
$f(x)=\frac{1}{A}\sum_{n\in\Upsilon}\left({\mathcal{E}}_{\psi}^{f}(\cdot,n)\star\psi_{n}\right)(x)=\frac{1}{A}\sum_{n\in\Upsilon}\int_{b\in{\mathbb{R}}^{N}}\langle f,T_{b}\psi_{n}\rangle T_{b}\psi_{n}(x)\mathrm{d}b.$
Now, we consider the symmetric wavelet filter bank $\{\chi_{n}\}_{n\in\Upsilon^{+}}$. The following theorem guarantees the exact reconstruction of a real-valued function $f$ from the continuous symmetric empirical wavelet transform ${\mathcal{E}}_{\chi}^{f}$ given in Equation 5.
@@
$f(x)=\sum_{n\in\Upsilon^{+}}\left({\mathcal{E}}_{\chi}^{f}(\cdot,n)\star\tilde{\chi}_{n}\right)(x)=\sum_{n\in\Upsilon^{+}}\int_{b\in{\mathbb{R}}^{N}}\langle f,T_{b}\chi_{n}\rangle T_{b}\widetilde{\chi}_{n}(x)\mathrm{d}b,$ (9)
where the set of dual symmetric empirical wavelets $\{\widetilde{\chi}_{n}\}_{n\in\Upsilon}$ is defined by, for every $n\in\Upsilon$ and $\xi\in{\mathbb{R}}^{N}$,
$\widehat{\widetilde{\chi}_{n}}(\xi)=\frac{\widehat{\chi}_{n}(\xi)}{\displaystyle\sum_{n\in\Upsilon^{+}}\left|\widehat{\chi}_{n}(\xi)\right|^{2}}.$
@@
$\sum_{n\in\Upsilon^{+}}\left|\widehat{\chi}_{n}(\xi)\right|^{2}$ $=\left|\widehat{\psi}_{0}(\xi)\right|^{2}+\frac{1}{2}\sum_{n\in\Upsilon^{+}\setminus\{0\}}\left|\widehat{\psi}_{n}(\xi)+\widehat{\psi}_{-n}(\xi)\right|^{2}$
@@
which ensures that the symmetric empirical wavelet filters $\widetilde{\chi}_{n}$ are well defined. Then, Equation 9 stems from the same computation as in the proof of Theorem 1. ∎
@@
then $\{T_{b}\chi_{n}\}_{(n,b)\in\Upsilon^{+}\times{\mathbb{R}}^{N}}$ is a continuous tight frame. Specifically, for every $x\in{\mathbb{R}}^{N}$,
$f(x)=\frac{1}{A+B}\sum_{n\in\Upsilon^{+}}\int_{b\in{\mathbb{R}}^{N}}\langle f,T_{b}\chi_{n}\rangle T_{b}\chi_{n}(x)\mathrm{d}b.$
@@
$\sum_{n\in\Upsilon^{+}}\left|\widehat{\chi}_{n}(\xi)\right|^{2}$ $=\frac{1}{2}\sum_{n\in\Upsilon^{+}}\left|\widehat{\psi}_{n}(\xi)\right|^{2}+\frac{1}{2}\sum_{n\in\Upsilon^{+}}\left|\widehat{\psi}_{-n}(\xi)\right|^{2}+\sum_{n\in\Upsilon^{+}\setminus\{0\}}\left|\widehat{\psi}_{n}(\xi)\right|\left|\widehat{\psi}_{-n}(\xi)\right|=A+B,$
and the result follows from the fact that $\widetilde{\chi}_{n}=\chi_{n}/(A+B)$. ∎
@@ 4.2 Discrete frames
In this section, we exhibit conditions to build discrete wavelet frames involved in the wavelet transforms of Definition 6. These conditions depend on the compactness of the support of the wavelet kernel’s Fourier transform $\widehat{\psi}$. The two underlying cases are examined separately.
@@ 4.2.1 Compactly supported $\boldsymbol{\widehat{\psi}}$
In this section, we consider that $\widehat{\psi}$ has a compact support. Excluding the supports of a partition $\{\Omega_{n}\}_{n\in\Upsilon}$ with non-compact closure, we can state sufficient and necessary conditions for which the systems
$\{T_{b_{n}k}\psi_{n}\}_{(n,k)\in\Upsilon_{\rm comp}\times{\mathbb{Z}}^{N}}\quad\textrm{ and }\quad\{T_{b_{n}k}\chi_{n}\}_{(n,k)\in\Upsilon_{\rm comp}^{+}\times{\mathbb{Z}}^{N}},$
@@
$\Upsilon_{\rm comp}=\{n\in\Upsilon\mid\overline{\Omega_{n}}\textrm{ is compact}\}\;\textrm{ and }\;\Upsilon_{\rm comp}^{+}=\{n\in\Upsilon^{+}\mid\overline{\Omega_{n}}\textrm{ is compact}\}.$
The following theorem first gives a sufficient and necessary condition to build a tight frame from $\{\psi_{n}\}_{n\in\Upsilon_{\rm comp}}$. It is a straight generalization of Theorems 4-7 of [10] to the $N$-dimensional case.
@@
Let us denote $\mathrm{L}_{\rm comp}^{2}({\mathbb{R}}^{N})=\{f\in\mathrm{L}^{2}({\mathbb{R}}^{N})\mid\mathrm{supp}\,\widehat{f}\subseteq\Gamma_{\rm comp}\}$ and $\Gamma_{\rm comp}=\underset{n\in\Upsilon_{\rm comp}}{\bigcup}\overline{\Omega_{n}}$. The system $\{T_{b_{n}k}\psi_{n}\}_{(n,k)\in\Upsilon_{\rm comp}\times{\mathbb{Z}}^{N}}$ is a Parseval frame for $\mathrm{L}_{\rm comp}^{2}({\mathbb{R}}^{N})$ if and only if, for $a.e.\,\xi\in{\mathbb{R}}^{N}$,
@@
$\mathcal{K}=\underset{n\in\Upsilon_{\rm comp}}{\bigcup}b_{n}^{-1}\mathbb{Z}^{N},\qquad\Upsilon_{\alpha}=\{n\in\Upsilon_{\rm comp}\mid b_{n}\alpha\in\mathbb{Z}^{N}\},$
@@
Let us denote $\mathcal{D}=\{f\in\mathrm{L}^{2}({\mathbb{R}}^{N})\mid\widehat{f}\in\mathrm{L}^{\infty}({\mathbb{R}}^{N})\textrm{ and }\mathrm{supp}\,\widehat{f}\subset\Gamma_{\rm comp}\}.$ Theorem 2.1 in [13] states, with $g_{p}=\psi_{n}$, $C_{p}=b_{n}$ and $\mathcal{P}=\Upsilon_{\rm comp}$, the desired equivalence under the condition that
$\forall f\in\mathcal{D},\;\sum_{n\in\Upsilon_{\rm comp}}\sum_{k\in\mathbb{Z}^{N}}\int_{\mathrm{supp}\,\widehat{f}}\left|\widehat{f}(\xi+b_{n}^{-1}k)\right|^{2}\frac{1}{|b_{n}|}\left|\widehat{\psi}_{n}(\xi)\right|^{2}\mathrm{d}\xi<\infty.$
@@
Similarly, the following theorem gives a sufficient and necessary condition to build a tight frame from the symmetric filter bank $\{\chi_{n}\}_{n\in\Upsilon_{\rm comp}^{+}}$.
Theorem 4 (Discrete symmetric Parseval frame).
Let us denote $\mathrm{L}_{\rm comp}^{2}({\mathbb{R}}^{N})=\{f\in\mathrm{L}^{2}({\mathbb{R}}^{N})\mid\mathrm{supp}\,\widehat{f}\subseteq\Gamma_{\rm comp}\}$ and $\Gamma_{\rm comp}=\underset{n\in\Upsilon_{\rm comp}}{\bigcup}\overline{\Omega_{n}}$. Then, the system $\{T_{b_{n}k}\chi_{n}\}_{(n,b)\in\Upsilon_{\rm comp}^{+}\times{\mathbb{Z}}^{N}}$ is a Parseval frame for $\mathrm{L}_{\rm comp}^{2}({\mathbb{R}}^{N})$ if and only if, for $a.e.\,\xi\in{\mathbb{R}}^{N}$,
$\sum_{n\in\Upsilon_{\alpha}^{+}}\frac{1}{|b_{n}|}\widehat{\chi}_{n}(\xi)\overline{\widehat{\chi}_{n}(\xi+\alpha)}=\delta_{\alpha,0},$ (10)
@@
$\mathcal{K}^{+}=\underset{n\in\Upsilon_{\rm comp}^{+}}{\bigcup}b_{n}^{-1}\mathbb{Z}^{N},\qquad\Upsilon_{\alpha}^{+}=\{n\in\Upsilon_{\rm comp}^{+}\mid b_{n}\alpha\in\mathbb{Z}^{N}\},$
@@
Let us denote $\mathcal{D}=\{f\in\mathrm{L}^{2}({\mathbb{R}}^{N})\mid\widehat{f}\in\mathrm{L}^{\infty}({\mathbb{R}}^{N})\textrm{ and }\mathrm{supp}\,\widehat{f}\subset\Gamma_{\rm comp}\}.$ Noticing that
$\Gamma_{\rm comp}=\underset{n\in\Upsilon_{\rm comp}^{+}}{\bigcup}\left(\overline{\Omega_{n}}\cup\overline{\Omega_{-n}}\right).$
Theorem 2.1 in [13] states (by taking $g_{p}=\chi_{n}$, $C_{p}=b_{n}$ and $\mathcal{P}=\Upsilon_{\rm comp}^{+}$) that the system $\{T_{b_{n}k}\chi_{n}\}_{(n,k)\in\Upsilon^{+}\times{\mathbb{Z}}^{N}}$ is a Parseval frame for $\mathrm{L}_{\rm comp}^{2}({\mathbb{R}}^{N})$ if and only if, for $a.e.\,\xi\in{\mathbb{R}}^{N}$,
$\sum_{n\in\Upsilon_{\alpha}^{+}}\frac{1}{|b_{n}|}\widehat{\chi}_{n}(\xi)\overline{\widehat{\chi}_{n}(\xi+\alpha)}=\delta_{\alpha,0},$
@@
$\forall f\in\mathcal{D},\;\sum_{n\in\Upsilon_{\rm comp}^{+}}\sum_{k\in\mathbb{Z}^{N}}\int_{\mathrm{supp}\,\widehat{f}}\left|\widehat{f}(\xi+b_{n}^{-1}k)\right|^{2}\frac{1}{|b_{n}|}\left|\widehat{\chi}_{n}(\xi)\right|^{2}\mathrm{d}\xi<\infty.$
This condition is given by the proof of Theorem 4 in [10], replacing $\mathbb{Z}$ by $\mathbb{Z}^{N}$ and $\psi_{n}$ by $\chi_{n}$. It results from the fact that $\mathrm{supp}\,\widehat{f}$ and $\mathrm{supp}\,\widehat{\chi}_{n}$ are compact and that there are finitely many $\mathrm{supp}\,\widehat{\psi}_{n}$ that intersect $\mathrm{supp}\,\widehat{f}$. This proves the equivalence. ∎